OpenStudy (anonymous):

) Determine the equation of the line passing through the points (1, 2) and (6, -4).

OpenStudy (anonymous):

y = mx + b can you find the slope m?

OpenStudy (anonymous):

Use Point-Slope form. y - y1 = m(x - x1) You need one point and the slope. So find slope(m): \[m = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ (-4) - 2 }{ 6 - 1 } = \frac{ -6 }{ 5 }\] Let's use point (1, 2). So: y - 2 = -6/5(x - 1) Distribute. Solve for y. y - 2 = -6/5x - 1.2 Add 2 to both sides. y = -6/5 + 0.8

OpenStudy (anonymous):

I meant: y = -6/5x + 0.8

OpenStudy (anonymous):

@jazy you did a good job but 0.8 is incorrect, do you want to retry

OpenStudy (anonymous):

Yes. (: Where did i make the first error?

OpenStudy (anonymous):

y - 2 = -6/5(x - 1) Distribute. Solve for y. y - 2 = -6/5x - 1.2 ---here

OpenStudy (anonymous):

It's supposed to be + 1.2, right? Or should it stay as a fraction?

OpenStudy (anonymous):

its + so now finish it

OpenStudy (anonymous):

Okay, my bad. y - 2 = -6/5x + 1.2 Add 2 to both sides. y = -6/5x + 3.2

OpenStudy (anonymous):

you got it, just be careful when multiplying ... other than that you are good at math

OpenStudy (anonymous):

Thanks, I'll keep that in mind! (:

OpenStudy (anonymous):

gg thanks aaalot

OpenStudy (anonymous):

i never new i should have changed it into a decimal